### v1

@@ 3.2.3 No Data Access for Vetted Researchers
The AIA does not contain a regulation comparable to Art. 40 DSA. This is surprising, as the regulation aims to ensure that relevant entities, such as digital innovation hubs, testing experimentation facilities and researchers, should be able to access and use high quality datasets within their respective fields of activities for the development and assessment of high-risk AI systems. [^65] Recital No. 45 (1) AIA. The European data rooms created by the AIA and the facilitated exchange of data should also provide trustful, accountable and non-discriminatory access to high quality data for the training, validation and testing of AI systems. [^66] Recital No. 45 (2) AIA.
@@ 4.1 Third-Party Audits by Researchers and Civil Society
[13, 74, 59, 45, 83] are some examples of algorithm audits in the United States that have uncovered problematic algorithmic behavior or algorithm impacts that could potentially cause harm. Regarding algorithmic audits in the EU, audits of the Austrian AMS algorithm [5], and Rotterdam’s welfare fraud algorithm by the Wire [17], come to mind. These audits were third-party audits, that is, audits by independent, external parties combining independent researchers, civil society, regulators, journalists, and other stakeholders. They have uncovered sociotechnical harms of algorithmic systems such as representational harms due to discriminatory choice.
@@ 4.1.2 Third-Party ADM and ML-Model Access
It appears to be just as challenging with ADMs and ML-driven products. Many of the products do not have access for researchers and regulatory audits. Therefore, mostly commercial APIs have to be used as in the Gender Shades case. If there is no API, the data can only be scraped, which is mostly illegal or prevented by other means. US Researchers are currently campaigning against the use of anti-hacking laws in the US that prevent audit studies by scraping. At the same time, according to Raji et al. [70], companies continue to seek to prevent audits by (1) paywalls, (2) prohibition laws through terms of services, and (3) structuring the product to obscure any clear set of test points. Raji et al. [70] describes it as an extraordinary effort for researchers, NGOs, investigative journalists to gain access and knowledge of ML and ADM systems.
@@ 4.1.3 Technical Methods and Access Levels
Document audits are often broad sets of information that can analysed in different ways [54] and ensure the possiblity for third-party impact evaluations or conformity assessments as intended by the AIA. However, only process-based compliance is feasible through document audits, there are no standards for most of these reports and they are impossible to verify if no other data access is provided [54].
Code audits entail direct access to a system’s code and involve the dynamic execution of code with various inputs to observe corresponding output behaviors [84]. While they are suitable for some audits, code audits might not be ideal for AI product audits because the need for training data. Furthermore, potential complexities in reverse engineering system behavior - especially for dynamically trained algorithms such as recommender systems - hamper code audits’ feasibility.
@@
What is accessed
Tools and possible evaluations
Feasible audit types
@@
Case study third-party audit(s)
Process access
Frameworks, checklists, model Cards, datasheets, method paper, technical reports
Checklists for explainability, robustness, fairness and privacy
Document audits (algorithmic risk assessment and algorithmic impact evaluations)
Audit Austrian AMS algorithm[5]
Model access (’black-box’)
Access to the predictor model $f(\cdot)$ using artificial data $x^{\ast}$
Feature relevance plots, partial dependency plots, adversarial attacks, adversarial fairness, statistical disclosure
Scraping, API, sock puppet audit and code Audit
Bloomberg stable diffusion audit[59], discrimination in online ad delivery, [83], auditing radicalization pathways on YouTube [74]
Input access
Access to the predictor model $f(\cdot)$ and training data $x$ to predict $f(x)$
Surrogate explanations, synthetic data, bias in outcome
Crowdsourced Audit686868just a fraction of training data $x_{p}\in x$, API audit with training data
Cooperative audit of pymetrics’ candidate screening software[93]
Outcome access (’grey-box’)
Access to the predictor model $f(\cdot)$, training data $x$ to predict $f(x)=\hat{y}$ and the actual outcome $y$
Accuracy of explanations, concept drift analysis, bias in opportunity, inversion attacks
API audit with training data and history of actual outcomes
ProRepublicas investigation of North- pointe’s recidivism risk scoring system [45]
Parameter control and learning goal access
Learning procedure and target $L$, parameters $\phi$, $L(f_{\phi}(x))$ can be retrained
@@
Stability of explanations, stability analysis, stability of bias metrics, functionality stealing, model complexity, stress-testing, trade-off of bias and loss metric, model extraction
Retraining of the actual model with different parameters
Audit of the Rotterdams’ welfare fraud ADM[17, 18]
white-box
Access to the whole architecture $f$ and all documents
@@
Documents and specific explanations, model selection and validation, model selection and development, model security evaluation
Experimental audits (interventions)
Accessible by first-parties and possibly second-parties
@@
Table 2: The Table presents leveled access modes from ’process access’ to ’white-box access’. We connected access modes with ’what is accessed’, tools and feasible audit types. The table must be read so that in each access mode possible tools, audit types and accesses are added to those that were in the level before. This may not be the case only for process access, since documents are not available in each access level. The table is inspired and derived from Koshiyama et al. [44], Koshiyama et al. [43] and Brown et al. [12]. As in the case of table 1, this table should provide only an overview. We derived the table via literature review and not via systematic methods. A systematic evaluation is probably not conductable as the field is still emerging. However, we want to show how data access and audit methods are connected and demonstrate the trade-off between access and revealed information.
@@
For example, Bloomberg’s stable diffusion audit [59], or in YouTube’s radicalization audit [74], the prediction model was accessed via the API. Thus, problematic behaviour like biases could already be shown in both cases. The original gender shades study as well operated within the constraints of accessing their audit targets via a commercial API, mirroring the actions of a user within these platforms and thus using a socket puppet [68]. Others had even more access. In the case of the cooperative audit of pymetrics’ candidate screening software the auditors had ’grey-box’ access, as the outcome is not feasible [^69] As it’s unknowable if someone would have performed well that didn’t get the job. for recruitment algorithms [^70] This audit was criticized as [95] claimed that the authors got funded by the company for reducing the scope of the audit. If this claim is justified, this would mean that the audit would not been conducted by independent external auditors, but dependent and thus internal auditors. The example again demonstrates the power of tech companies.[94]. In the case of ProRepublicas investigation of the recidivism risk scoring system they had access to the model, the training data and the actual outcome [45]. In the case of Rotterdam’s welfare fraud audit, the auditors had parameter control and learning goal access [17]. Model access alone has shown its effectiveness through certain case study audits. Audits, including Bloomberg’s stable diffusion audit [59], examinations of discrimination in online ad delivery [83], and the audit of radicalization pathways on YouTube [74], have achieved auditing outcomes with limited access, albeit frequently encountering challenges due to legal constraints. Especially in the case of gender shades, where after the audit there was an improvement in several facial recognition tools in terms of functionality for black women [68]. This proofs limited ’black-box’ access for researchers can contribute to mitigating harm.
@@
While other levels of access might provide more comprehensive insights into a system, the prevalence of reverse-engineering methods underscores the significant influence of third-party audits. Admittedly, a system-tailored approach is necessary, given the nuances involved, and model-agnostic audits, such as partial dependency plots and feature relevance plots, present challenges [35]. Nonetheless, it is evident that third-party audits, even with ’black-box’ access, have managed to provide insights when combined with evaluations of the sociotechnical framework. Particularly because they contribute to a system for which the depth of knowledge acquired corresponds to the access level, audits with even minimal access can generate momentum for obtaining greater access. In some cases, this may involve collaborating with other auditors and stakeholders. This concept is further explored in the next chapter, “AI Audit Ecosystem.” It is worth noting that substantial examination by sock puppets or API audits is often feasible only for products that possess open APIs.
@@ 4.2 An AI Audit Ecosystem
In order to ensure oversight and compliance, Raji et al. [70] and Albert [3] have proposed establishing an AI audit ecosystem that includes internal and external auditors. Sambasivan et al. [75] have called for the establishment of an ecosystem for accountability in India that includes various stakeholders and empowers oppressed communities. Mökander et al. [56], meanwhile, has argued that the AIA de facto proposes to establish an AI audit ecosystem. However, none of these authors have described what constitutes an audit ecosystem or why we need one. While following this line of reasoning about the importance of including diverse auditors, we define an audit ecosystem and argue that the AIA does not establish an AI audit ecosystem.
@@
The audit ecosystem is needed to ensure compliance and oversight throughout the AI lifecycle. We argue that for such an ecosystem, third-party audits by researchers and civil society are a vital component to fulfill the characteristics of the ecosystem and thus also provide both oversight and compliance through first-party audits. Novelli et al. [60] have stated that ”[o]verseers may act at different levels that may overlap, e.g., an internal audit is compatible with judicial review accountability.” However, we argue that overseers should act at different levels, as (1) internal audits are prone to audit-washing as well as false assurances, and they may only ensure compliance (see section 4.3.1) and are (2) not adaptive to new technologies (see section 4.3.3). Additionally we argue that (3) NGOs and researchers possess valuable experience in incorporating a holistic perspective that encompasses socioecological-technical systems. They are able to draw from a background in conducting impact assessments and human rights considerations (see section 4.4.1) and (4) tend to include affected communities more than companies as they have other audience and incentives (see section 4.4.2).
@@ 4.3 Why Are Internal Audits Insufficient?
Furthermore, the overreliance on internal audits could lead to unverifiable assertions that the AI systems has passed ethical and legal standards, leading to even more harm and less oversight[32]. There is an ongoing discussion within research communities on how to translate ethics into practice. The discussion evolves around which fairness metrics should be used while as fairness remains a contested term [55]. General standards and norms for audits [19] are therefore absent. This leads to an environment conducive to the phenomenon of “audit-washing”. A firm’s business goals may not always align with harm-reduction through course-correction or ethical design of their products. Developers and deployers may ignore audit recommendations that threaten their business interests [58].
Empirical research substantiate these concerns. [52] have discovered that instructing developers to incorporate ethical codes yielded no notable change in their established practices. Moreover, studies involving AI development companies [88] and startup-like environments[87] have shown a disconnect between acknowledging the significance of ethics and integrating them into AI practices, indicating a substantial gap between ethical research and practical implementation and auditing.
Such absence of standards and norms in corporate environments could lead to using tools and metrics that play down the risks of ones product [39]. Statistically, through p-hacking or data dredging, and more generally through fairness gerrymandering, it may always be possible to suggest that your system is fair [8]. Issues concerning independence of second-party auditors were discussed frequently in financial audits. The opinion paper on this issue of Bazerman et al. [9] highlights the the emerging bias towards the client in contractual relationship audits. In addition to intentional false assurances of the absence of harm, audits may be performed incorrectly or issues may go unaddressed. That is what happened in the Gender Shades audit. It revealed that the U.S. National Institute of Technology (NIST) failed to measure demographic differences in their regular Facial Recognition Vendor’s Test [70].
@@ 4.3.3 Non-adaptivity to New AI Technologies Through Hard Standards
In an audit system of hard standards, ’gaming the system’ can occur from companies trying to exploit those standards and find loopholes. At the same time, new technologies in particular can create standardization gaps. To create an adaptive and dynamic audit ecosystem informal audits without hard standards and processes could help to find harm.
Social accountability in the public sector has demonstrated that strict process-based evaluations often slow down government action and suppress creativity. Therefore performance-based evaluations in audits are preferable to process-based ones. The role of civil society can be crucial in that regard, as it functions as a vigilant watchdog over both procedural integrity and performance outcomes. The involvement of civil society can introduce innovative performance metrics to supplement existing measures [1].
This becomes particularly evident if there is no external oversight from third parties capable of reproducing these evaluations. Novel AI products and technologies could necessitate the development of new evaluation methods. For example, in the aforementioned Gender Shades case or in recent examples of LLMs and stable diffusion. Although the relevant legislation and standards do not yet exist, it is difficult to imagine that if evaluation standards had existed, they would have worked for these type of AI systems. Developing standards for evaluating technology will take a long time, which is why it is necessary that third parties with limited access to the model engage in evaluating and understanding emerging systems. This happened with the stable diffusion audit by Bloomberg[59]  [^74] Reporters used stable diffusion to generate 5,100 images of people using a simple recurring prompt iterating through different categories. Through data-labeling and with ChatGPT, which researchers extensively evaluated and which showed biases [72].
4.4 Why Should We Include Research and Civil Society in the AI Audit Ecosystem?

### v3

@@ 3.2.3 No Data Access for Vetted Researchers
The AIA does not contain a regulation comparable to Art. 40 DSA. This is surprising, as the regulation aims to ensure that relevant entities, such as digital innovation hubs, testing experimentation facilities and researchers, should be able to access and use high quality datasets within their respective fields of activities for the development and assessment of high-risk AI systems. [^65] Rec. (68) AIA. The European data space created by the AIA and the facilitated exchange of data should also provide trustful, accountable and non-discriminatory access to high quality data for the training, validation and testing of AI systems. [^66] Recital No. 45 AIA.
@@ 4.1 Third-Party Audits by Researchers and Civil Society
Buolamwini and Gebru [14], Ribeiro et al. [80], Nicoletti and Bass [65], Larson et al. [49], Sweeney [89] are some examples of algorithm audits in the United States that have uncovered problematic algorithmic behavior or algorithm impacts that could potentially cause harm. Regarding algorithmic audits in the EU, audits of the Austrian AMS algorithm [5], and Rotterdam’s welfare fraud algorithm by the Wire [20], come to mind. These audits were third-party audits, that is, audits by independent, external parties combining independent researchers, civil society, regulators, journalists, and other stakeholders. They were responsible for uncovering deeply-rooted sociotechnical harms in algorithmic systems related mainly to representational harms due to discriminatory design choices. As a result, they showed the extent of the potential harms brought by AI systems’ biases.
@@ 4.1.2 Third-Party ADM and ML-Model Access
It appears to be just as challenging with ADMs and ML-driven products. Many of the products do not have access for researchers and regulatory audits. Therefore, mostly commercial APIs have to be used as in the Gender Shades case [14]. If there is no API, the data can only be scraped, which is mostly illegal or prevented by other means. US Researchers are currently campaigning against the use of anti-hacking laws in the US that prevent audit studies by scraping. At the same time, according to Raji et al. [76], companies continue to seek to prevent audits by (1) paywalls, (2) prohibition laws through terms of services, and (3) structuring the product to obscure any clear set of test points. Raji et al. [76] describes it as an extraordinary effort for researchers, NGOs, investigative journalists to gain access and knowledge of ML and ADM systems.
@@ 4.1.3 Technical Methods and Access Levels
Document audits are often broad sets of information that can be analysed in different ways [60] and ensure the possiblity for third-party impact evaluations or conformity assessments as intended by the AIA. However, only process-based compliance is feasible through document audits, there are no standards for most of these reports and they are impossible to verify if no other data access is provided [60].
Code audits entail direct access to a system’s code and involve the dynamic execution of code with various inputs to observe corresponding output behaviors [90]. While they are suitable for some audits, code audits might not be ideal for AI product audits because of the need for training data. Furthermore, potential complexities in reverse engineering system behavior - especially for dynamically trained algorithms such as recommender systems - hamper code audits’ feasibility.
@@
What is accessed Tools and possible evaluations Feasible audit types Case study third-party audit(s)
Process access Frameworks, checklists, model Cards, datasheets, method paper, technical reports Checklists for explainability, robustness, fairness and privacy Document audits (algorithmic risk assessment and algorithmic impact evaluations) Audit Austrian AMS algorithm[5]
Model access (’black-box’) Access to the predictor model $f(\cdot)$ using artificial data $x^{\ast}$ Feature relevance plots, partial dependency plots, adversarial attacks, adversarial fairness, statistical disclosure Scraping, API, sock puppet audit and code Audit Bloomberg stable diffusion audit[65], discrimination in online ad delivery, [89], auditing radicalization pathways on YouTube [80]
Input access Access to the predictor model $f(\cdot)$ and training data $x$ to predict $f(x)$ Surrogate explanations, synthetic data, bias in outcome Crowdsourced Audit686868just a fraction of training data $x_{p}\in x$, API audit with training data Cooperative audit of pymetrics’ candidate screening software[100]
Outcome access (’grey-box’) Access to the predictor model $f(\cdot)$, training data $x$ to predict $f(x)=\hat{y}$ and the actual outcome $y$ Accuracy of explanations, concept drift analysis, bias in opportunity, inversion attacks API audit with training data and history of actual outcomes ProRepublicas investigation of North- pointe’s recidivism risk scoring system [49]
@@
Parameter control and learning goal access Learning procedure and target $L$, parameters $\phi$, $L(f_{\phi}(x))$ can be retrained Stability of explanations, stability analysis, stability of bias metrics, functionality stealing, model complexity, stress-testing, trade-off of bias and loss metric, model extraction Retraining of the actual model with different parameters Audit of the Rotterdams’ welfare fraud ADM[20, 21]
@@
White-box access Access to the whole architecture $f$ and all documents Documents and specific explanations, model selection and validation, model selection and development, model security evaluation Experimental audits (interventions) Accessible by first-parties and possibly second-parties
@@
Table 2: The Table presents leveled access modes from ’process access’ to ’white-box access’. We connected access modes with ’what is accessed’, tools and feasible audit types. The table must be read so that in each access mode possible tools, audit types and accesses are added to those that were in the level before. This may not be the case only for process access, since documents are not available in each access level. The table is inspired and derived from Koshiyama et al. [48], Koshiyama et al. [47] and Brown et al. [13]. As in the case of table 1, this table should provide only an overview. We derived the table via literature review and not via systematic methods. A systematic evaluation may not be feasible as the field is still emerging. However, we aim to illustrate the connection between data access and audit methods and demonstrate the trade-off between access and revealed information.
@@
For example, Bloomberg’s stable diffusion audit [65], or in YouTube’s radicalization audit [80], the prediction model was accessed via the API. Thus, problematic behaviour like biases could already be shown in both cases. The original Gender Shades study as well operated within the constraints of accessing their audit targets via a commercial API, mirroring the actions of a user within these platforms and thus using a socket puppet [74]. Others had even more access. In the case of the cooperative audit of pymetrics’ candidate screening software the auditors had ‘grey-box’ access, as the outcome is not feasible [^69] As it’s unknowable if someone would have performed well that did not get the job. for recruitment algorithms [^70] This audit was criticized as [103] claimed that the authors got funded by the company for reducing the scope of the audit. If this claim is justified, this would mean that the audit would not been conducted by independent external auditors, but dependent and thus internal auditors. The example again demonstrates the power of tech companies.[101]. In the case of ProRepublicas investigation of the recidivism risk scoring system they had access to the model, the training data and the actual outcome [49]. In the case of Rotterdam’s welfare fraud audit, the auditors had parameter control and learning goal access [20]. Model access alone has shown its effectiveness through certain case study audits. Audits, including Bloomberg’s stable diffusion audit [65], examinations of discrimination in online ad delivery [89], and the audit of radicalization pathways on YouTube [80], have achieved auditing outcomes with limited access, albeit frequently encountering challenges due to legal constraints. Especially in the case of Gender Shades, where after the audit there was an improvement in several facial recognition tools in terms of functionality for black women [74]. This proofs limited ’black-box’ access for researchers can contribute to mitigating harm.
@@
While other levels of access might provide more comprehensive insights into a system, the prevalence of reverse-engineering methods underscores the significant influence of third-party audits. Admittedly, a system-tailored approach is necessary, given the nuances involved, and model-agnostic audits, such as partial dependency plots and feature relevance plots, present challenges [39]. Nonetheless, it is evident that third-party audits, even with ’black-box’ access, have managed to provide insights when combined with evaluations of the sociotechnical framework. Particularly because they contribute to a system for which the depth of knowledge acquired corresponds to the access level, audits with even minimal access can generate momentum for obtaining greater access. In some cases, this may involve collaborating with other auditors and stakeholders. This concept is further explored in the next section (section 4.2) It is worth noting that substantial examination by sock puppets or API audits is often feasible only for products that possess open APIs.
@@ 4.2 An AI Audit Ecosystem
In order to ensure oversight and compliance, Raji et al. [76] and Albert [3] have proposed establishing an AI audit ecosystem that includes internal and external auditors. Sambasivan et al. [81] have called for the establishment of an ecosystem for accountability in India that includes various stakeholders and empowers oppressed communities. Mökander et al. [62], meanwhile, have argued that the AIA de facto proposes to establish an AI audit ecosystem. However, none of these authors have described what constitutes an audit ecosystem or why we need one. While following this line of reasoning about the importance of including diverse auditors, we define an audit ecosystem and argue that the AIA does not establish an AI audit ecosystem.
@@
The audit ecosystem is needed to ensure compliance and oversight throughout the AI lifecycle. We argue that for such an ecosystem, third-party audits by researchers and civil society are a vital component to fulfill the characteristics of the ecosystem and thus also provide both oversight and compliance through first-party audits. Novelli et al. [66] have stated that “[o]verseers may act at different levels that may overlap, e.g., an internal audit is compatible with judicial review accountability.” However, we argue that overseers should act at different levels, as (1) internal audits are prone to audit-washing as well as false assurances, and they may only ensure compliance (see section 4.3.1) and are (2) not adaptive to new technologies (see section 4.3.3). Additionally we argue that (3) NGOs and researchers possess valuable experience in incorporating a holistic perspective that encompasses socioecological-technical systems. They are able to draw from a background in conducting impact assessments and human rights considerations (see section 4.4.1) and (4) tend to include affected communities more than companies as they have other audience and incentives (see section 4.4.2).
@@ 4.3 Why Are Internal Audits Insufficient?
Furthermore, the overreliance on internal audits could lead to unverifiable assertions that the AI systems has passed ethical and legal standards, leading to even more harm and less oversight [36]. There is an ongoing discussion within research communities on how to translate ethics into practice. The discussion evolves around which fairness metrics should be used as fairness remains a contested term [61]. General standards and norms for audits [22] are therefore absent. This leads to an environment conducive to the phenomenon of “audit-washing”. A firm’s business goals may not always align with harm-reduction through course-correction or ethical design of their products. Developers and deployers may ignore audit recommendations that threaten their business interests [64].
Empirical research substantiate these concerns. McNamara et al. [58] have discovered that instructing developers to incorporate ethical codes yielded no notable change in their established practices. Moreover, studies involving AI development companies [95] and startup-like environments[94] have shown a disconnect between acknowledging the significance of ethics and integrating them into AI practices, indicating a substantial gap between ethical research and practical implementation and auditing.
Such absence of standards and norms in corporate environments could lead to using tools and metrics that play down the risks of ones own product [43]. Statistically, through p-hacking or data dredging, and more generally through fairness gerrymandering, it may always be possible to suggest that your system is fair [8]. Issues concerning independence of second-party auditors were discussed frequently in financial audits. The opinion paper on this issue of Bazerman et al. [9] highlights the the emerging bias towards the client in contractual relationship audits. In addition to intentional false assurances of the absence of harm, audits may be performed incorrectly or issues may go unaddressed. That is what happened in the Gender Shades audit [14]. It revealed first assessment and benchmarks by the U.S. National Institute of Technology (NIST) failed to measure demographic differences in their regular Facial Recognition Vendor’s Test [76].
@@ 4.3.3 Non-adaptivity to New AI Technologies Through Hard Standards
Internal audits and hard standards have several drawbacks. First, social accountability in the public sector has demonstrated that strict process-based evaluations often slow down government action and suppress creativity. Therefore performance-based evaluations in audits are preferable to process-based ones. The role of civil society can be crucial in that regard, as it functions as a vigilant watchdog over both procedural integrity and performance outcomes. The involvement of civil society can introduce innovative performance metrics to supplement existing measures [1].
Especially the AI standardization process, as currently envisaged under the AIA, has its challenges with respect to that. In this standardization process, harmonized standards are adopted following a standardization request (mandate) from the Commission to the European Standardization Organizations (CEN, CENELEC, building the CEN-CENELEC JTC21 technical joint committee) for the application of Union harmonization legislation to harmonize AI standards [57].
This process was criticized several times [57, 102, 52]. First, the process itself is non-participatory in nature [102]. Second, the imposition of hard standards on ethical questions that are inherently context-sensitive and dynamic with new technologies is problematic [52].
Third, in an audit ecosystem of hard standards – like in the AIA case – ‘gaming the system’ can occur from companies trying to exploit those standards and find loopholes. At the same time, new technologies in particular can create standardization gaps. To create an adaptive and dynamic audit ecosystem informal audits without hard standards and processes could help to expose potential harm.
This becomes particularly evident if there is no external oversight from third parties capable of reproducing these evaluations. Novel AI products and technologies could necessitate the development of new evaluation methods. New ethical questions arise, which may have to include ethical trade-offs navigated by local and diverse stakeholder inclusion. For example, in the aforementioned Gender Shades case [14] or in recent examples of LLMs and stable diffusion [65]. Although the relevant legislation and standards do not yet exist, it is difficult to imagine that if evaluation standards had existed, they would have worked for these type of AI systems. Developing standards for evaluating technology will take a long time, which is why it is necessary that third parties with limited access to the model engage in evaluating and understanding emerging systems. This happened with the stable diffusion audit by Bloomberg[65]  [^74] Reporters used stable diffusion to generate 5,100 images of people using a simple recurring prompt iterating through different categories. Through data-labeling and with ChatGPT, which researchers extensively evaluated and which showed biases [78].
4.4 Why Should We Include Researchers and Civil Society in the AI Audit Ecosystem?